Let Y be a random variable with PDF

F(y) = {(1/2)(1-y) -1≤ y ≤ 1

{ 0 elsewhere

a) Find the density function of X = 1 - 2Y

b) Find the density function of U = Y^2
Answer #1

Given, pdf f(y) = (1-y)/2 ; -1\leqy\leq 1

To find the density function of the respective transformations, the distribution function method is employed.

(a) x 21-24 P(x=xc)2 P(1-24 Ex) zP(Y 21-x) 21- P(y=1-2) 21- Fj (1-x) polt godt 2-4 (1-3) = 1 =(-1*) When y 2-1 0 23 Yz1 xzod  Ф) из уа Plus u) = P(y²=u) 2 Pl- Jusys vu) 2 Plysru)- Ply=-5w) 2 Fy (vw) - Fytou) Pdf of u, hW datat (rw) + frattua) na Ilva)